# Kai Shu, Lu Cheng, Huan Liu

### v5

@@ 4.2.2 Rule Augmentation
Rule-augmented prompting involves embedding a set of principles, references, and evaluation rubrics directly within the prompt for the judge LLM. This approach is commonly employed in LLM-based evaluations, where judge LLMs are guided to assess specific aspects Li et al. (2024e); Bai et al. (2023a); Yu et al. (2024d); Qian et al. (2024) and provided with detailed rubrics Gao et al. (2023b); Kim et al. ; Wang et al. (2024f); Murugadoss et al. (2024); Li et al. (2024l, h); Chen et al. (2024c) to ensure a fair comparison. A distinct approach is seen in Liu et al. (2024b), where the judge LLM is prompted to generate its own scoring criteria through in-context learning on a set of few-shot examples. In alignment with LLM-as-a-judge, Bai et al. (2022) first propose to introduce a list of principles (e.g., helpfulness, harmlessness, honesty) for the judge LLM to compare the two candidates more precisely and directionally. Following them, subsequent works Lee et al. (2023, 2024); Guo et al. (2024); Sun et al. (2024a); Beigi et al. (2024a) enhance this principle-driven prompting by incorporating more detailed explanations for each aspect of the principle or rubric. Besides, both Li and Qiu (2023) and Li et al. (2024c) propose to prompt LLMs to retrieve appropriate demonstrations/ knowledge triples based on the candidates’ helpfulness in solving specific problems. To obtain diverse responses from LLMs, Lahoti et al. (2023) prompt multiple LLMs to judge the diversity of each candidate and select the most diverse one for further polishing. Zhang et al. (2024g) propose RevisEval, which leverages the self-correction capabilities of LLMs to adaptively revise the response, then treat the revised text as the principles for the subsequent evaluation. Recently,some works investigate the reliability of LLM-as-a-personalized-judge Dong et al. (2024); Seraj et al. (2024); Kim and Baek (2024), providing persona as a part of principles for LLMs to make personalized judgments.
@@ 5.1.2 Reasoning Tasks
The reasoning capability of LLMs can be evaluated through their intermediate thinking process and final answers on specific reasoning tasks Mondorf and Plank (2024); Bernard et al. (2024). Recently, LLM-as-a-judge has been used to evaluate the logical progression, depth, and coherence of the model’s intermediate reasoning paths. For mathematical reasoning tasks, Xia et al. (2024) introduce an automatic evaluation framework using a judge LLM specifically designed to assess the quality of reasoning steps in problem-solving processes. LLM-as-a-judge can also be applied to more complex reasoning tasks like temporal reasoning, where models need to understand the relationships of different events over time. Fatemi et al. (2024) build synthetic datasets specifically tailored to evaluate LLMs’ temporal reasoning abilities across varied scenarios, testing their proficiency in reasoning with sequences, causality, and dependencies among temporally ordered events.
@@ 5.1.3 Emerging Tasks
Recently, we have seen more customized utilization of LLM-as-a-judge to evaluate emerging tasks such as code understanding Zhao et al. (2024a); Quan et al. (2025); Yu et al. (2024c); Feng et al. (2024a); Lu et al. (2024a); Biagiola et al. (2024), legal knowledge Fei et al. (2023), game development Isaza-Giraldo et al. (2024), ocean science Bi et al. (2023), healthcare conversations Wang et al. (2024m), debating judgment Liang et al. (2024a), retrival-augmented generation Dhole et al. (2024), cross ability Zhong et al. (2024), HCI Li et al. (2024j); Chai et al. (2025); Goren et al. (2024), role-playing Zhou et al. (2024b), RAG Jin et al. (2024), speech synthesis Park et al. (2024b), counterspeech generation Bennie et al. (2025b, a) and more. This trend reflects the growing adaptability of LLMs-as-judges in evaluating diverse and specialized domains.
@@ 5.2.1 Larger Models as Judges
One intuitive idea for adopting LLM-as-a-judge in alignment tuning is to leverage the feedback from larger, more powerful LLMs to guide smaller, less capable models. Bai et al. (2022) first propose employing the AI’s feedback to build a harmless AI assistant. They train the reward model using synthetic preference data based on the preference of a pre-trained language model. Building on this, Lee et al. (2023) discover that the RLAIF method can achieve comparable performance with RLHF even when the LLM judge is not strong enough. They also introduced DIRECT-RLAIF, which directly employs an off-the-shelf LLM as the judge model to mitigate reward staleness in reward models. To avoid reward hacking in alignment, Sun et al. (2024a) devise an instructable reward model trained on synthetic preference data. It enables humans to perform RL-time interventions to better align the target policy with human values. Apart from the abovementioned studies, Guo et al. (2024) introduce online AI feedback (OAIF), directly utilizing the preference signals from an annotation model to train the target model. There are also works that utilize multi-agent cooperation for better judgment in alignment tuning. Arif et al. (2024) and Sengupta et al. (2024) construct a synthetic preference optimization dataset using multi-agent workflows and adopt LLMs as judges with diverse prompting strategies and pipelines. Similarly, Li et al. (2024i) employ multiple LLMs to debate with each other, iteratively improving response quality, while creating a judge LLM to select preferred responses for enhanced instruction tuning. To align the generated code with human preference, Weyssow et al. (2024) introduce CodeUltraFeedback, a preference coding dataset constructed using the LLM-as-a-Judge methodology. This synthetic dataset is later used to fine-tune and align small code LLMs using SFT and DPO. Recently, Wang et al. (2024e) propose BPO, using GPT-4 as the judge and constructing synthetic pairwise feedback for knowledge depth and breadth balance in the alignment process.
@@ 5.2.2 Self-Judging
Another line of work aims to utilize the preference signal from the same LLM for self-improving. Yuan et al. (2024e) first propose the concept of a self-rewarding LLM, where pairwise data is constructed by having the LLM itself serve as the judge. Following them, Wu et al. (2024a) introduce meta-rewarding that judges the LLMs’ judge and uses the feedback to refine their judging skills. Their LLM-as-a-meta-judge approach significantly enhances the model’s capacity to evaluate and follow instructions. To improve synthetic data quality, Pace et al. (2024) combine the Best-of-N and Worst-of-N sampling strategies and introduce the West-of-N approach. Lee et al. (2024) devise Judge augmented Supervised Fine-Tuning (JSFT) to train a single model to act as both a policy and a judge. To fully utilize this judge model, they also propose the Self-Rejection by Tournament method to select the best response in the inference time. Unlike the above approaches, which use the LLM as a judge for pairwise data construction, Tong et al. (2024) apply LLM-as-a-judge as a self-filtering method to ensure the quality of synthetic data pairs in alignment tasks for reasoning. To reduce the computation overhead in pairwise judgment, Zhai et al. (2024) propose a ranked pairing method for self-preferring language models, which accelerates the comparison process by measuring the strength of each response against a baseline. Liu et al. (2024c) introduce meta-ranking, enabling weaker LLMs Sun et al. (2024b) to serve as reliable judges and provide trustworthy feedback. They also apply their meta-ranking method in post-SFT training, combining it with Kahneman-Tversky Optimization (KTO) for improved alignment. To enhance the quality of the synthetic instruction tuning data, Liang et al. (2024c) introduce an iterative self-enhancement paradigm (I-SHEEP). During training, they adopt LLM-as-a-judge to score the synthetic responses and set a threshold to collect high-quality query-response pairs for the subsequent training iteration. Recently, Yasunaga et al. (2024) propose to combine LLM-as-a-judge and data synthesis and successfully build human-aligned LLMs with a few annotation.
@@ 5.4.1 Reasoning Path Selection
Additionally, there are also many works that focus on step-level reasoning path selection, leveraging the judge LLM as the process reward model (PRM) to evaluate state scores. Creswell et al. (2023) decompose the reasoning process into Selection and Inference. In the selection step, they leverage the LLM itself to judge and evaluate each potential reasoning trace, selecting the appropriate one for the following inference step. Xie et al. (2024b) propose the Kwai-STaR framework, which transforms LLMs into state-transition reasoners to judge and select the best reasoning state for themselves in mathematical reasoning. Lightman et al. (2023) train a LLM as the PRM to conduct inference-time supervision and perform best-of-N sampling strategy during reasoning stage. Following them, Setlur et al. (2024) further propose process advantage verifiers (PAVs), generating rewards based on the changes in the likelihood of producing correct responses in the future. Other works simulate advanced cognitive structures as the reasoning process. Hao et al. (2023) employ LLMs as a world model to simulate the state of the environment and perform Monte Carlo Tree Search (MCTS) to improve performance on tasks requiring deliberate path selection. Furthermore, Besta et al. (2024) takes the outputs generated by LLMs as an arbitrary graph. An LLM thought is modeled as a vertex, while an edge is a dependency between thoughts. This framework enables a systematic judge of coherence and logical reasoning for each reasoning state. Yao et al. (2023a) propose Tree-of-Thoughts (ToT), where each thought serves as an intermediate step toward problem-solving. It decomposes reasoning into steps, self-evaluates and judges progress at each state, and uses search algorithms with LMs to judge thought paths through lookahead and backtracking. Besides, there are also studies train critique-based LLM judges Ankner et al. (2024); Yu et al. (2024b); Wang et al. (2024d); Lan et al. ; Xie et al. (2024b); Yao et al. (2024); Yuksel and Sawaf (2024) which provide fine-grained verbal feedback to boost the reasoning process.
@@
DHP Wang et al. (2024i) Various Tasks 400 Human Discernment Score General Performance
@@ 5.4.2 Reasoning with External Tools
Yao et al. (2023b) first propose to use LLMs in an interleaved manner to generate reasoning traces and task-specific actions. Reasoning traces help the model to judge and update action plans, while actions enable it to interact with external sources. Further, Auto-GPT was introduced by Yang et al. (2023) to deliver more accurate information with LLM-as-a-judge for tool-using. By equipping with a range of external complex tools, LLMs become more versatile and capable, improving planning performance by judging and reasoning which tools to use. Sha et al. (2023) explore LLMs’ potential in reasoning and judging, employing them as decision-making components for complex autonomous driving scenarios that require human commonsense understanding. Zhou et al. (2024c) utilize a self-discovery process where LLMs perform judgment based on the given query and select the most feasible reasoning structure for the following inference stage.
@@ 6.1 General Performance
  • JUDGE-BENCH Tan et al. (2024a), DHP Wang et al. (2024i), RewardBench Lambert et al. (2024), SOS-BENCH Penfever et al. (2024) and JuStRank Gera et al. (2024), which operate at larger scales, utilizing metrics like Cohen’s kappa, Discernment Score, and normalized accuracy to benchmark general LLM performance.
@@ 7.1 Bias & Vulnerability
LLM judges are also highly susceptible to adversarial manipulations. Techniques like JudgeDeceiver highlight the risks posed by optimization-based prompt injection attacks, where carefully crafted adversarial sequences can manipulate LLM judgments to favor specific responses Shi et al. (2024a); BENCHMARK ; Banerjee et al. (2024). Similarly, universal adversarial phrases can drastically inflate scores in absolute scoring paradigms, revealing vulnerabilities in zero-shot assessment setups Liusie et al. (2023); Raina et al. (2024); Doddapaneni et al. (2024). These manipulations raise concerns about the reliability of LLM judges in high-stakes scenarios such as leaderboards and academic or legal assessments Shi et al. (2024a); Raina et al. (2024); Zhang et al. (2024e); Jiang et al. (2024).
@@
Future Direction. One promising direction for future research is to integrate Retrieval-Augmented Generation (RAG) frameworks into LLM evaluation processes Chen et al. (2024e). By combining generative and retrieval capabilities, these frameworks can reduce biases such as self-preference and factuality issues by grounding evaluations in external, verifiable data sources. Another promising avenue is the use of bias-aware datasets, like OFFSETBIAS, to systematically address inherent biases in LLM-as-a-judge systems Park et al. (2024a). Incorporating such datasets into training pipelines allows LLMs to better distinguish superficial qualities from substantive correctness, thereby enhancing fairness and reliability. Exploring fine-tuned LLMs as scalable judges, exemplified by the JudgeLM framework, represents another intriguing direction Zhu et al. (2023). Techniques like swap augmentation and reference support embedded in this framework can systematically mitigate biases, improve evaluation consistency, and extend the applicability of LLM-based judgments to open-ended tasks. Additionally, advancing zero-shot comparative assessment frameworks offers significant promise Liusie et al. (2023). These frameworks can refine pairwise comparison techniques and implement debiasing strategies, improving fairness and reliability across diverse evaluation domains without the need for extensive prompt engineering or fine-tuning. Finally, methods like JudgeDeceiver-resistant calibration and adversarial phrase detection strategies need further exploration to secure LLM-as-a-judge frameworks from attacks Shi et al. (2024a); Raina et al. (2024); Zhou et al. (2024a); Li et al. (2024d).

### v6

@@ 4.2.2 Rule Augmentation
Rule-augmented prompting involves embedding a set of principles, references, and evaluation rubrics directly within the prompt for the judge LLM. This approach is commonly employed in LLM-based evaluations, where judge LLMs are guided to assess specific aspects Li et al. (2024e); Bai et al. (2023a); Yu et al. (2024d); Qian et al. (2024) and provided with detailed rubrics Gao et al. (2023b); Kim et al. ; Wang et al. (2024g); Murugadoss et al. (2024); Li et al. (2024l, h); Chen et al. (2024c) to ensure a fair comparison. A distinct approach is seen in Liu et al. (2024d), where the judge LLM is prompted to generate its own scoring criteria through in-context learning on a set of few-shot examples. In alignment with LLM-as-a-judge, Bai et al. (2022) first propose to introduce a list of principles (e.g., helpfulness, harmlessness, honesty) for the judge LLM to compare the two candidates more precisely and directionally. Following them, subsequent works Lee et al. (2023, 2024); Guo et al. (2024); Sun et al. (2024a); Beigi et al. (2024a) enhance this principle-driven prompting by incorporating more detailed explanations for each aspect of the principle or rubric. Besides, both Li and Qiu (2023) and Li et al. (2024c) propose to prompt LLMs to retrieve appropriate demonstrations/ knowledge triples based on the candidates’ helpfulness in solving specific problems. To obtain diverse responses from LLMs, Lahoti et al. (2023) prompt multiple LLMs to judge the diversity of each candidate and select the most diverse one for further polishing. Zhang et al. (2024g) propose RevisEval, which leverages the self-correction capabilities of LLMs to adaptively revise the response, then treat the revised text as the principles for the subsequent evaluation. Recently,some works investigate the reliability of LLM-as-a-personalized-judge Dong et al. (2024); Seraj et al. (2024); Kim and Baek (2024), providing persona as a part of principles for LLMs to make personalized judgments.
@@ 5.1.2 Reasoning Tasks
The reasoning capability of LLMs can be evaluated through their intermediate thinking process and final answers on specific reasoning tasks Mondorf and Plank (2024); Bernard et al. (2024); Wang et al. (2024c). Recently, LLM-as-a-judge has been used to evaluate the logical progression, depth, and coherence of the model’s intermediate reasoning paths. For mathematical reasoning tasks, Xia et al. (2024) introduce an automatic evaluation framework using a judge LLM specifically designed to assess the quality of reasoning steps in problem-solving processes. LLM-as-a-judge can also be applied to more complex reasoning tasks like temporal reasoning, where models need to understand the relationships of different events over time. Fatemi et al. (2024) build synthetic datasets specifically tailored to evaluate LLMs’ temporal reasoning abilities across varied scenarios, testing their proficiency in reasoning with sequences, causality, and dependencies among temporally ordered events.
@@ 5.1.3 Emerging Tasks
Recently, we have seen more customized utilization of LLM-as-a-judge to evaluate emerging tasks such as code understanding Zhao et al. (2024a); Quan et al. (2025); Yu et al. (2024c); Feng et al. (2024a); Lu et al. (2024a); Biagiola et al. (2024), legal knowledge Fei et al. (2023), game development Isaza-Giraldo et al. (2024), ocean science Bi et al. (2023), healthcare conversations Wang et al. (2024n), debating judgment Liang et al. (2024a), retrival-augmented generation Dhole et al. (2024), cross ability Zhong et al. (2024), HCI Li et al. (2024j); Chai et al. (2025); Goren et al. (2024), role-playing Zhou et al. (2024c), RAG Jin et al. (2024), speech synthesis Park et al. (2024b), counterspeech generation Bennie et al. (2025b, a) and more. This trend reflects the growing adaptability of LLMs-as-judges in evaluating diverse and specialized domains.
@@ 5.2.1 Larger Models as Judges
One intuitive idea for adopting LLM-as-a-judge in alignment tuning is to leverage the feedback from larger, more powerful LLMs to guide smaller, less capable models. Bai et al. (2022) first propose employing the AI’s feedback to build a harmless AI assistant. They train the reward model using synthetic preference data based on the preference of a pre-trained language model. Building on this, Lee et al. (2023) discover that the RLAIF method can achieve comparable performance with RLHF even when the LLM judge is not strong enough. They also introduced DIRECT-RLAIF, which directly employs an off-the-shelf LLM as the judge model to mitigate reward staleness in reward models. To avoid reward hacking in alignment, Sun et al. (2024a) devise an instructable reward model trained on synthetic preference data. It enables humans to perform RL-time interventions to better align the target policy with human values. Apart from the abovementioned studies, Guo et al. (2024) introduce online AI feedback (OAIF), directly utilizing the preference signals from an annotation model to train the target model. There are also works that utilize multi-agent cooperation for better judgment in alignment tuning. Arif et al. (2024) and Sengupta et al. (2024) construct a synthetic preference optimization dataset using multi-agent workflows and adopt LLMs as judges with diverse prompting strategies and pipelines. Similarly, Li et al. (2024i) employ multiple LLMs to debate with each other, iteratively improving response quality, while creating a judge LLM to select preferred responses for enhanced instruction tuning. To align the generated code with human preference, Weyssow et al. (2024) introduce CodeUltraFeedback, a preference coding dataset constructed using the LLM-as-a-Judge methodology. This synthetic dataset is later used to fine-tune and align small code LLMs using SFT and DPO. Recently, Wang et al. (2024f) propose BPO, using GPT-4 as the judge and constructing synthetic pairwise feedback for knowledge depth and breadth balance in the alignment process.
@@ 5.2.2 Self-Judging
Another line of work aims to utilize the preference signal from the same LLM for self-improving. Yuan et al. (2024e) first propose the concept of a self-rewarding LLM, where pairwise data is constructed by having the LLM itself serve as the judge. Following them, Wu et al. (2024a) introduce meta-rewarding that judges the LLMs’ judge and uses the feedback to refine their judging skills. Their LLM-as-a-meta-judge approach significantly enhances the model’s capacity to evaluate and follow instructions. To improve synthetic data quality, Pace et al. (2024) combine the Best-of-N and Worst-of-N sampling strategies and introduce the West-of-N approach. Lee et al. (2024) devise Judge augmented Supervised Fine-Tuning (JSFT) to train a single model to act as both a policy and a judge. To fully utilize this judge model, they also propose the Self-Rejection by Tournament method to select the best response in the inference time. Unlike the above approaches, which use the LLM as a judge for pairwise data construction, Tong et al. (2024) apply LLM-as-a-judge as a self-filtering method to ensure the quality of synthetic data pairs in alignment tasks for reasoning. To reduce the computation overhead in pairwise judgment, Zhai et al. (2024) propose a ranked pairing method for self-preferring language models, which accelerates the comparison process by measuring the strength of each response against a baseline. Liu et al. (2024e) introduce meta-ranking, enabling weaker LLMs Sun et al. (2024b) to serve as reliable judges and provide trustworthy feedback. They also apply their meta-ranking method in post-SFT training, combining it with Kahneman-Tversky Optimization (KTO) for improved alignment. To enhance the quality of the synthetic instruction tuning data, Liang et al. (2024c) introduce an iterative self-enhancement paradigm (I-SHEEP). During training, they adopt LLM-as-a-judge to score the synthetic responses and set a threshold to collect high-quality query-response pairs for the subsequent training iteration. Recently, Yasunaga et al. (2024) propose to combine LLM-as-a-judge and data synthesis and successfully build human-aligned LLMs with a few annotation.
@@ 5.4.1 Reasoning Path Selection
Additionally, there are also many works that focus on step-level reasoning path selection, leveraging the judge LLM as the process reward model (PRM) to evaluate state scores. Creswell et al. (2023) decompose the reasoning process into Selection and Inference. In the selection step, they leverage the LLM itself to judge and evaluate each potential reasoning trace, selecting the appropriate one for the following inference step. Xie et al. (2024b) propose the Kwai-STaR framework, which transforms LLMs into state-transition reasoners to judge and select the best reasoning state for themselves in mathematical reasoning. Lightman et al. (2023) train a LLM as the PRM to conduct inference-time supervision and perform best-of-N sampling strategy during reasoning stage. Following them, Setlur et al. (2024) further propose process advantage verifiers (PAVs), generating rewards based on the changes in the likelihood of producing correct responses in the future. Other works simulate advanced cognitive structures as the reasoning process. Hao et al. (2023) employ LLMs as a world model to simulate the state of the environment and perform Monte Carlo Tree Search (MCTS) to improve performance on tasks requiring deliberate path selection. Furthermore, Besta et al. (2024) takes the outputs generated by LLMs as an arbitrary graph. An LLM thought is modeled as a vertex, while an edge is a dependency between thoughts. This framework enables a systematic judge of coherence and logical reasoning for each reasoning state. Yao et al. (2023a) propose Tree-of-Thoughts (ToT), where each thought serves as an intermediate step toward problem-solving. It decomposes reasoning into steps, self-evaluates and judges progress at each state, and uses search algorithms with LMs to judge thought paths through lookahead and backtracking. Besides, there are also studies train critique-based LLM judges Ankner et al. (2024); Yu et al. (2024b); Wang et al. (2024e); Lan et al. ; Xie et al. (2024b); Yao et al. (2024); Yuksel and Sawaf (2024) which provide fine-grained verbal feedback to boost the reasoning process.
@@
DHP Wang et al. (2024j) Various Tasks 400 Human Discernment Score General Performance
@@ 5.4.2 Reasoning with External Tools
Yao et al. (2023b) first propose to use LLMs in an interleaved manner to generate reasoning traces and task-specific actions. Reasoning traces help the model to judge and update action plans, while actions enable it to interact with external sources. Further, Auto-GPT was introduced by Yang et al. (2023) to deliver more accurate information with LLM-as-a-judge for tool-using. By equipping with a range of external complex tools, LLMs become more versatile and capable, improving planning performance by judging and reasoning which tools to use. Sha et al. (2023) explore LLMs’ potential in reasoning and judging, employing them as decision-making components for complex autonomous driving scenarios that require human commonsense understanding. Zhou et al. (2024d) utilize a self-discovery process where LLMs perform judgment based on the given query and select the most feasible reasoning structure for the following inference stage.
@@ 6.1 General Performance
  • JUDGE-BENCH Tan et al. (2024a), DHP Wang et al. (2024j), RewardBench Lambert et al. (2024), SOS-BENCH Penfever et al. (2024) and JuStRank Gera et al. (2024), which operate at larger scales, utilizing metrics like Cohen’s kappa, Discernment Score, and normalized accuracy to benchmark general LLM performance.
@@ 7.1 Bias & Vulnerability
LLM judges are also highly susceptible to adversarial manipulations. Techniques like JudgeDeceiver highlight the risks posed by optimization-based prompt injection attacks, where carefully crafted adversarial sequences can manipulate LLM judgments to favor specific responses Shi et al. (2024a); BENCHMARK ; Banerjee et al. (2024); Zhou et al. (2024a); Liu et al. (2024b). Similarly, universal adversarial phrases can drastically inflate scores in absolute scoring paradigms, revealing vulnerabilities in zero-shot assessment setups Liusie et al. (2023); Raina et al. (2024); Doddapaneni et al. (2024). These manipulations raise concerns about the reliability of LLM judges in high-stakes scenarios such as leaderboards and academic or legal assessments Shi et al. (2024a); Raina et al. (2024); Zhang et al. (2024e); Jiang et al. (2024).
@@
Future Direction. One promising direction for future research is to integrate Retrieval-Augmented Generation (RAG) frameworks into LLM evaluation processes Chen et al. (2024e). By combining generative and retrieval capabilities, these frameworks can reduce biases such as self-preference and factuality issues by grounding evaluations in external, verifiable data sources. Another promising avenue is the use of bias-aware datasets, like OFFSETBIAS, to systematically address inherent biases in LLM-as-a-judge systems Park et al. (2024a). Incorporating such datasets into training pipelines allows LLMs to better distinguish superficial qualities from substantive correctness, thereby enhancing fairness and reliability. Exploring fine-tuned LLMs as scalable judges, exemplified by the JudgeLM framework, represents another intriguing direction Zhu et al. (2023). Techniques like swap augmentation and reference support embedded in this framework can systematically mitigate biases, improve evaluation consistency, and extend the applicability of LLM-based judgments to open-ended tasks. Additionally, advancing zero-shot comparative assessment frameworks offers significant promise Liusie et al. (2023). These frameworks can refine pairwise comparison techniques and implement debiasing strategies, improving fairness and reliability across diverse evaluation domains without the need for extensive prompt engineering or fine-tuning. Finally, methods like JudgeDeceiver-resistant calibration and adversarial phrase detection strategies need further exploration to secure LLM-as-a-judge frameworks from attacks Shi et al. (2024a); Raina et al. (2024); Zhou et al. (2024b); Li et al. (2024d); Liu et al. (2024c).